# disposition

### v1

@@ 4 Nonlinear Cone Separation Conditions
$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)$ =$=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
@@
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then,
$x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma\quad\text{ for all }t\geq 0$
and
$\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty\quad\text{ for }t\to\infty.$
Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
@@
On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$. By the boundedness of $\Omega_{2}$ there is $\lambda>0$ such that $||x||\leq\lambda$ for all $x\in\Omega_{2}$. Consequently, we have
$||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0\quad\text{for }t_{n}\to\infty.$
This means that $0\in\operatorname{cl}\,\Omega_{1}\subseteq\operatorname{cl}_{w}\,\Omega_{1}$. Noting that the compact (respectively, weakly compact) set $\Omega_{1}$ is closed (respectively, weakly closed), we get a contradiction.
@@
To later prove a Bishop-Phelps type scalarization result (namely Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
@@ 5 Bishop-Phelps Type Scalarization in Vector Optimization
Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$. Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as
$\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)\quad\text{for all }y\in Y.$
The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
@@ 5.1 Basic scalarization results
Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
@@
$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid. Now assume that $\operatorname{int}\,K\neq\emptyset$. If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so the function $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$). The proof of the reverse implication is similar to $1^{\circ}$.
@@
Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
@@
$2^{\circ}$. For the case $\operatorname{int}\,K=\emptyset$ the statement is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Here the latter set is contained in ${\mathrm{W}Eff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
@@
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}-3^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed, then the assertions $1^{\circ}-4^{\circ}$ are equivalent to:
@@
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
@@
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}$ and $2^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then the assertions $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
@@
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
    $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
@@
  • Gasimov [17, Th. 2] (see also Guo and Zhang [31, Th. 4.5]) stated
    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$
    $\Longrightarrow\{x\in\Omega$ $\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$
    with the cone-valued map $\mathbb{A}$ considered in Benson proper efficiency.
@@
Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence
$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K).$
If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g. if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23)
$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K).$
If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence
$\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K).$
@@ 5.2 Scalarization results for the concept of weak efficiency
Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
$(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$
hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
@@ 5.3 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
Under our assumptions, by Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$
Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
$(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$
hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
@@
Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result is a direct consequence of Theorem 5.25.
In the remaining part of the subsection, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.

### v2

@@ 4 Nonlinear Cone Separation Conditions
$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
@@
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$. Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
@@
On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$ for $n\to\infty$. By the boundedness of $\Omega_{2}$ there is $\lambda>0$ such that $||x||\leq\lambda$ for all $x\in\Omega_{2}$. Consequently, we have $||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0$ for $n\to\infty$. This means that $0\in\operatorname{cl}\,\Omega_{1}\subseteq\operatorname{cl}_{w}\,\Omega_{1}$. Noting that the compact (respectively, weakly compact) set $\Omega_{1}$ is closed (respectively, weakly closed), we get a contradiction.
@@
To later prove a Bishop-Phelps type scalarization result (Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
@@ 5 Bishop-Phelps Type Scalarization in Vector Optimization
Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$. Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as $\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)$ for all $y\in Y$. The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
@@ 5.1 Basic scalarization results
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
@@
$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid. Assume that $\operatorname{int}\,K\neq\emptyset$. If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$). The proof of the reverse implication is similar to $1^{\circ}$.
@@
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
@@
$2^{\circ}$. For the case $\operatorname{int}\,K=\emptyset$ the statement is clear. Assume that $\operatorname{int}\,K\neq\emptyset$. Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Here the latter set is contained in ${\mathrm{W}Eff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
@@
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed, then $1^{\circ}-4^{\circ}$ are equivalent to:
@@
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
@@
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
@@
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
@@
  • Gasimov [17, Th. 2] (see also Guo and Zhang [31, Th. 4.5]) stated that $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
@@
Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$. If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g. if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23) $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
@@ 5.2 Scalarization results for the concept of weak efficiency
Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
@@ 5.3 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
By Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$. Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
@@
Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result follows from Theorem 5.25.
In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.